OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - pi/3), what should be used for Xmin and Xmax?

OpenStudy (anonymous):

you have an infinite number of choices

OpenStudy (anonymous):

What would you start with/recommend? @satellite73

OpenStudy (anonymous):

set \[2(x-\frac{\pi}{3})=0\] and \[2(x-\frac{\pi}{3})=4\pi\] and solve both for \(x\) would be one way to do it

OpenStudy (anonymous):

Thank you @satellite73

OpenStudy (anonymous):

yw